The electric potential at points in an xy plane is givenby V = (3.0 V/m2)x2 -(3.8V/m2)y2. What are(a) the magnitude of the electric field at thepoint (4.3 m, 1.7 m) and (b) the angle that thefield there makes with the positive x direction.
Answer #1
We know the electricfield in a direction is partial derivative of potential w.r.t thatdirection
so Ex= δV/δx = δ( (3.0V/m2)x2 -(3.8V/m2)y2)/ δx= 6x V/m

at x=4.3

Ex= 4.3*6 = 25.8N/C
Ey= δV/δy = δ( (3.0V/m2)x2 -(3.8V/m2)y2)/ δy= 7.6 y V/m

at y =1.7
Ey = 1.7*7.6 = 12.92N/C
Magnitude = sqrt (12.92 2 + 25.8 2 ) = 28.9 N/c

Angle with x axis =cos -1(Ex/Enet) = cos -1 (25.8/28.9)